OpenStudy (anonymous):

I NEED HELP !!!!!!!!! <3 (only people willing to help me)

OpenStudy (saifoo.khan):

with?

OpenStudy (bahrom7893):

okay here: l = w+7

OpenStudy (bahrom7893):

A = 30 A = 30 = l * w = (w+7)w w^2 + 7w = 30

OpenStudy (anonymous):

i dont believe you :/

OpenStudy (bahrom7893):

W = 3 L = w+7 = 10 Fine if u dont want to believe me..

OpenStudy (anonymous):

your a meany .. if i get it wrong i fail my test and im retained .. but ima go with it i guess :(

OpenStudy (bahrom7893):

How am I a meany?

OpenStudy (zarkon):

then maybe you should actually learn the material and stop cheating.

OpenStudy (anonymous):

do i put 3,10 ?